# Accelerating AllReduce with a Persistent Straggler

### v1

@@ Appendix G End-to-End Training Simulations at Scale
We use the FlexNet simulator used by prior work [49]—an augmented version of FlexFlow [51, 88]—to determine 3D parallelism schemes for transformer training at even larger scales. Specifically, FlexNet outputs the task graph of compute and communication tasks per training iteration. We simulate training the BERT-large transformer model [89] as we vary both the number of GPUs and the straggler delay. We provide end-to-end training iteration time reported by FlexNet and using the analytical network simulator to predict the communication time (for both collective and point-to-point communication) as an input. Fig. 10 shows end-to-end results assuming $\alpha{=}3\mu{s}$ (based on  [33]) while Fig. 11 shows these results assuming $\alpha{=}0.7\mu{s}$ (based on  [20]). In general, StragglAR outperforms baselines as the number of GPUs scales when there is even marginal delay of 0.5 ms or higher. When there is no delay, StragglAR suffers because none of the ReduceScatter precondition is overlapped with the straggler delay. In Fig. 10, StragglAR either performs more poorly than or evenly with RHD (depending on the straggler delay) for $n{=}64$ because the assumption of $\alpha$ is higher (so, algorithms like RHD that minimize the number of rounds are preferred when $n$ is large). In contrast, Fig. 11 assumes a lower value of $\alpha$ that is reported by prior published work [20], thereby enabling StragglAR’s bandwidth efficiency to shine even at large values of $n$. The exact buffer sizes and compute and communication kernels are directly outputted by FlexNet, and we evaluated with batch sizes of 16, 32, 64, and 128 on 8, 16, 32, and 64 GPUs, respectively. These batch sizes are based on typical finetuning use cases that opt for smaller batch sizes on smaller datasets [85, 86, 87].

### v2

@@ Appendix I End-to-End Training Simulations at Scale
We use the FlexNet simulator used by prior work [Wang et al., 2023a]—an augmented version of FlexFlow [Jia et al., 2019, FlexFlow, ]—to determine 3D parallelism schemes for transformer training at even larger scales. Specifically, FlexNet outputs the task graph of compute and communication tasks per training iteration. We simulate training the BERT-large transformer model [Kenton and Toutanova, 2019] as we vary both the number of GPUs and the straggler delay. We provide end-to-end training iteration time reported by FlexNet and using the analytical network simulator to predict the communication time (for both collective and point-to-point communication) as an input. Fig. 12 shows end-to-end results assuming $\alpha{=}3\mu{s}$ (based on [Microway, 2024]) while Fig. 13 shows these results assuming $\alpha{=}0.7\mu{s}$ (based on [Shah et al., 2023]). In general, StragglAR outperforms baselines as the number of GPUs scales when there is even marginal delay of 0.5 ms or higher. When there is no delay, StragglAR suffers because none of the ReduceScatter precondition is overlapped with the straggler delay. In Fig. 12, StragglAR either performs more poorly than or evenly with RHD (depending on the straggler delay) for n=64n{=}64 because the assumption of $\alpha$ is higher (so, algorithms like RHD that minimize the number of rounds are preferred when $n$ is large). In contrast, Fig. 13 assumes a lower value of $\alpha$ that is reported by prior published work [Shah et al., 2023], thereby enabling StragglAR’s bandwidth efficiency to shine even at large values of $n$. The exact buffer sizes and compute and communication kernels are directly outputted by FlexNet, and we evaluated with batch sizes of 16, 32, 64, and 128 on 8, 16, 32, and 64 GPUs, respectively. These batch sizes are based on typical fine-tuning use cases that opt for smaller batch sizes on smaller datasets [Chung et al., 2024, Liu et al., 2019, Touvron et al., 2023].